OpenStudy (anonymous):

Find the 6 trig functions P(-2square root of 5,-square root of 5)

OpenStudy (jdoe0001):

so you have a point, two coordinates \(\Large {\begin{matrix} P\left(-\frac{2}{5}, \sqrt{5}\right)\\ &-\frac{2}{5}&\sqrt{5}\\ &x&y\\ &a& b \end{matrix}}\)  to find "c", use the pythagorean theorem once you have all 3 guys, "a", "b", and "c" then get all functions :)

OpenStudy (jdoe0001):

hmm is negative "y", missed that there, so \(\Large \begin{matrix} P\left(-\frac{2}{5}, \sqrt{5}\right)\\ &-\frac{2}{5}&-\sqrt{5}\\ &x&y\\ &a& b \end{matrix}\)

OpenStudy (jdoe0001):

woohoo, typos, anyhow, \(\Large \begin{matrix} P\left(-\frac{2}{5}, -\sqrt{5}\right)\\ &-\frac{2}{5}&-\sqrt{5}\\ &x&y\\ &a& b \end{matrix}\)

OpenStudy (anonymous):

But how did you get -2over 5

OpenStudy (anonymous):

\[\Large \begin{matrix} P\left(-2\sqrt{5}, -\sqrt{5}\right)\\ \end{matrix}\]

OpenStudy (anonymous):

Yes @cinar but how do I find the third number and the six trig functions

OpenStudy (anonymous):